VELO

#11071 Sensor with errors + lower pseudo efficiency

Started at 2018-05-03 09:46. Severity: Unknown. 1 comments

2018-05-03 10:42 Victor Coco Two sensors show a much higher rate of tracks with one used sensor from a module than the reference. It seems to correspond, in the Velo monitoring plots, too sensors with high number of error, the same sensors have lower pseudo-efficinecy vs sensor ID. I checked few runs in the passed, it seems to have been there for some time although less in the past (closer to reference) See logbook for plots

TT

#11192 Bins at values 42.9 42.9 42.9 43.0 are empty

Started at 2018-05-19 23:47. Severity: Report. 1 comments

2018-05-19 23:47 Aleksandr Dolmatov TT alarm: Dead links compared to reference. algorithm CheckDeadBins:Bins at values 42.9 42.9 42.9 43.0 are empty. The expert said the alarm would stay for the whole night. If the problem concerns only these 4 bins, then nothing to worry about.

#11191 TT Cluster: hit map. See empty bins in TTbV. (from Presenter page /Shift/TT: Clusters)

Started at 2018-05-19 14:06. Severity: Report. 1 comments

2018-05-19 14:06 Mengzhen Wang Around 6 bins are empty

#11186 See alarm in "Number of clusters vs TELL1 links": Dead links compared to reference

Started at 2018-05-19 08:31. Severity: Report. 1 comments

2018-05-19 08:31 Mengzhen Wang Informed the ST piquet.

#11183 TT error banks: Fraction of events containing an error bank is around 100% for bin 42. (from Presenter page /Shift/TT: Error banks)

Started at 2018-05-19 07:34. Severity: Report. 1 comments

2018-05-19 07:34 Mengzhen Wang This is not fixed by changing run

OT

#11155 See a lot of error message in the OT:Errors histograms from 17:25 - 17:35 (from Presenter page /Shift/OT: Errors)

Started at 2018-05-14 17:54. Severity: Report. 1 comments

2018-05-14 17:54 Mengzhen Wang See a lot of error message in the OT:Errors histograms from 17:25 - 17:35 (from Presenter page /Shift/OT: Errors)

MUON

#11034 Hole in M1A region R4

Started at 2018-04-12 11:32. Severity: Minor. 2 comments

2018-04-12 11:37 Oleg Maev Small hole in M1A R4, chamber 9C. This chamber lost signals from two FEBs. Impact on efficiency is very minor.
2018-04-29 07:26 Vladimir Macko Hole still observed, for plot provided for a future reference in the logbook

L0

#11188 SPD L0CALO: Status: error rate of Spd multiplicity reaches 100%. error rate of photon reaches ~0.01. (from Presenter page /Shift/L0CALO: Status)

Started at 2018-05-19 10:44. Severity: Report. 1 comments

2018-05-19 10:44 Mengzhen Wang SPD L0CALO: Status: error rate of Spd multiplicity reaches 100%. error rate of photon reaches ~0.01. (from Presenter page /Shift/L0CALO: Status)

#11178 L0Calo error rate too large (from Presenter page /Shift/L0CALO: Status)

Started at 2018-05-18 10:28. Severity: Report. 1 comments

2018-05-18 10:28 Mengzhen Wang L0Calo error rate is visible for Electron, Photon and Hadron.

#11154 some peaks appear in "Number of Bad Channels" histogram (from Presenter page /Shift/L0MUON: Status)

Started at 2018-05-12 15:55. Severity: Report. 1 comments

2018-05-12 15:55 Mengzhen Wang Some peaks are seen in the histogram of "Number of Bad Channels". They disappear very soon and none of the peaks last more than 5min, but the value of the peak is quite high, which can reach 2.5.

HLT

#11189 Error banks seen by Hlt1 during 11:30 - 11:40 (from Presenter page /Shift/HLT: 2- Error Banks seen by HLT1)

Started at 2018-05-19 11:47. Severity: Report. 1 comments

2018-05-19 11:47 Mengzhen Wang Error banks seen by Hlt1 during 11:30 - 11:40 (from Presenter page /Shift/HLT: 2- Error Banks seen by HLT1)

#11185 L0 rate goes to zero around 07:45 (from Presenter page /Shift/HLT: 1- L0 and HLT1 Rates)

Started at 2018-05-19 08:03. Severity: Report. 1 comments

2018-05-19 08:03 Mengzhen Wang BTW, it seems that the x-axis for the L0 rate plot in the presenter is 1 hour smaller than what it should be...

TRACK

#11190 Track: z closest to z-axis (from Presenter page /Shift/TRACK: Track properties and PV alignment)

Started at 2018-05-19 12:59. Severity: Report. 1 comments

2018-05-19 12:59 Mengzhen Wang The data and reference fit not very well in the tail in the left.

#11174 Velo alignment: PV left-right delta x mean value > 0.01mm (from Presenter page /Shift/TRACK: Velo Alignment)

Started at 2018-05-17 11:34. Severity: Report. 1 comments

2018-05-17 11:34 Mengzhen Wang The mean value of the delta x distribution is around 0.011 mm. Compared to the reference, the shift is around 0.005mm.